OpenStudy (anonymous):

How do you graph y+3^x?

OpenStudy (anonymous):

With a pencil and paper!

OpenStudy (anonymous):

Do u mean y= -3^x? or do u mean f(x,y) = y+3^x

OpenStudy (anonymous):

Sorry, couldn't resist.

OpenStudy (anonymous):

Wish I could like that. Lol

OpenStudy (anonymous):

And the second one

OpenStudy (anonymous):

3d graph.

OpenStudy (anonymous):

I don't have a calculator so I can't enter it in to see what it looks like

OpenStudy (anonymous):

Click the link.

OpenStudy (anonymous):

And this is the first one (2D) http://www.wolframalpha.com/input/?i=y%3D+-3^x

OpenStudy (anonymous):

THANK YOU:) :")

OpenStudy (anonymous):

Ur welcome.